# Elementary commutator method for the Dirac equation with long-range perturbations

### v1

@@ 2.5 Absence of super-exponentially decaying eigenstates
$\kappa\|\bar{\chi}_{a}^{1/2}\mathrm{e}^{\kappa(f-2^{b+2})}\phi\|^{2}$ $\leq C\|\chi_{b-1,b+1}\phi\|^{2}.$ (2.36)
Now we assume $\bar{\chi}_{b+2}\phi\neq 0$ as a function on $\mathbb{R}^{d}$. Then the left-hand side of (2.36) grows exponentially as $\kappa\to\infty$ whereas the right-hand side remains bounded. This is a contradiction, and thus $\bar{\chi}_{b+2}\phi=0$. Then by the unique continuation property [10] we conclude that $\phi=0$ as a function on $\mathbb{R}^{d}$. We are done. ∎
@@ 3.1.2 Commutator estimate
for short, and proceed similarly to (2.14) and (2.15) as
@@
$\mathop{\mathrm{Im}}\bigl(A_{\Theta}(H-z)\bigr)\geq(c_{1}-q_{1})\theta^{\prime}-\mathop{\mathrm{Re}}(\gamma(H-z))-C_{10}Q,$ (3.8)
@@ Acknowledgements
S.A. is supported by Forefront Physics and Mathematics Program to Drive Transformation (FoPM), a World-leading Innovative Graduate Study (WINGS) Program, the University of Tokyo. K.I. is partially supported by JSPS KAKENHI Grant Number 23K03163. A part of this work was done while K.I. visited Aarhus University. He would like to express his gratitude for kind hospitality by Arne Jensen and Erik Skibsted there.

### v2

@@ 2.4 Absence of super-exponentially decaying eigenstates
$\bigl\|\bar{\chi}_{a}^{1/2}\mathrm{e}^{\kappa(f-2^{a+2})}\phi\bigr\|^{2}\leq C\|\chi_{a-1,a+1}\phi\|^{2},$ (2.31)
with $\bar{\chi}_{a}$ being from (2.7). Now we assume $\bar{\chi}_{a+2}\phi\not\equiv 0$ on $\mathbb{R}^{d}$. Then the left-hand side of (2.31) grows exponentially as $\kappa\to\infty$, whereas the right-hand side remains bounded. This is a contradiction, and thus $\bar{\chi}_{b+2}\phi\equiv 0$. Then by the unique continuation property [10] we conclude that $\phi\equiv 0$ on $\mathbb{R}^{d}$. We are done. ∎
@@ 3.1.2 Commutator estimate
for short, and proceed similarly to (2.15), (2.17) and (2.18) as
@@
$\mathop{\mathrm{Im}}\bigl(A_{\Theta}(H-z)\bigr)\geq(c_{1}-q_{1})\theta^{\prime}-\mathop{\mathrm{Re}}(\gamma(H-z))-C_{9}Q,$ (3.8)
@@ Acknowledgements
S.A. is supported by Forefront Physics and Mathematics Program to Drive Transformation (FoPM), a World-leading Innovative Graduate Study (WINGS) Program, the University of Tokyo. K.I. is partially supported by JSPS KAKENHI Grant Number 23K03163. A part of this work was done while K.I. visited Aarhus University. He would like to express his gratitude for kind hospitality by Arne Jensen and Erik Skibsted there. The authors would like to sincerely thank Kyohei Itakura for pointing out a serious error in the previous version of the paper.